From GlueXWiki
Jump to: navigation, search

May 9, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
    • Spare package status
  2. Update on the corrosion test at MePhi (Lubomir)
  3. Installation (Lubomir)
  4. Engineering (Bill)
    • Cooling system test
  5. Electronics (Chris, Nick)
  6. Full electronics tests FDC E-log (Beni)
  7. Other


Participants: Bill, Chris, Nick, Simon, Beni, and Lubomir.


- All the wire frames are deadened now.

- Spare package installation: Two cells were installed last week. The bottom cell (wire frames #2) was tested and works fine, but the top one (wire frame #3) showed different problems: no signal on the first wire (suspect wire is broken between the pad and epoxy), point noise source on one wire, glowing discharge appearing from time to time at >2kV. The discharge can be identified by a sudden increase of the currents on both sense and field wires up to 1-10 uA that doesn't go away unless the voltage is decreased. Instead of trying to fix wire frame #3 we decided to replace it with wire frame #6. While testing #6 before the installation, two broken wires were found and then fixed by Casey in two days: one of the wires was giving somehow signals on the first day. Finally this morning we installed the second cell with wire frame #6 that will be tested on Monday.

Update on the corrosion test at MePhi

- Vlad sent us update about his latest results: the test at 65degC had six, only EPDM, samples (no chance to see some effects with Viton at this temperature) and they failed in between 30 to 54 hours. In the first test (at 100degC) the two EPDM samples were found with high resistivity in 8 and 24 hours, but the latter sample wasn't tested in between. Therefore Vlad decided to repeat the test at 100degC only with EPDM samples (six) but this time checking the samples more often in the first 24 hours.


- The draft of the installation document is ready and linked above. Beni wrote the section about the DAQ and trigger requirements for the FDC test outside of the magnet. One can use the document when writing the installation procedures; there are enough details for that. Bill (probably also with Tim) will look at the paper and will give us his/their recommendations/changes/remarks.

- In particular we discussed that frames that will support the top and bottom scintillator planes (each with two paddles) needed for the testing. Beni will give Bill information about the paddles and Bill will sketch the frames.


- Casey replaced the manifolds on package #3 with aluminum ones and then the cooling system was tested with 100 psi pressure from the CO2 bottle: no leakage found with the sniffer.

- Bill about the cooling system test at 126: yesterday we tested the system with Fluorinert, now with aluminum manifolds no leakages were found except at some points on the copper tubing where it was machined (drilled to increase the inner diameter). However, the water chiller appeared to have a problem when working with the SMC chiller. The SMC chiller stops the flow of the cold water from time to time when it's not needed. In case of no water flow, the pump of the water chiller starts heating and the water temperature increases up to 45degC. To fix the problem Bill wants to bypass the SMC chiller if it blocks the water and is working on that. Another option will be to replace the pump with another type, but will cost more.

Full electronics tests

- Beni: Cody and Bryan still working on the firmware/software for fADC125. Beni will get more cables from Nick for the LV system in 126.

- After the meeting Eugene proposed to install the third package in 126, install all the fADC125s needed, and connect them with cables. Even if we can't read now the new fADCs we can look at the noise of the old fADC125s that we can read. The problem is that the modules may have to be moved back to Cody to install the firmware. Beni will talk to Cody and Brian about this option, but in any case we plan the move the third package next week.